A 49.4 V emf  is placed across a series combination of 3 resistors of size 10 ?, 38.767 ?

and 37 ?.

At what rate is heat generated in the 38.767 ? resistor?
Answer #1

i = emf/(R1 + R2 + R3) = 49.4/(10+38.767+37) = 0.57598 A


P = R i^2 = 38.767 * (0.57598*0.57598) = 12.9 W
Answer #2

first calculate current in all of the resistors -

I = V/ R.eq = 49.4/(85.767) = 0.575 A R.eq = eqvivalent resistance

now heat generated in a resistor = (I^2 )* R

so H = (0.575^2)*38.767 = 12.86 Joule